OpenStudy (emilyjones284):

Let z=1- square root of 3i. Find z^7

zepdrix (zepdrix):

\[\Large\bf\sf z\quad=\quad 1-\sqrt3\mathcal i\]Factoring out a 2 from each term,\[\Large\bf\sf z\quad=\quad 2\left(\frac{1}{2}-\frac{\sqrt3}{2}\mathcal i\right)\]We can write it in terms of sine and cosine from here. Following along so far?

OpenStudy (emilyjones284):

Yeah I got it so far

zepdrix (zepdrix):

The imaginary component is negative, real is positive.. so that puts ussssss in theeeeeee uhhhhhh... 4th quadrant. So our angle is 5pi/3, I think.

zepdrix (zepdrix):

In exponential form we have:\[\Large\bf\sf z\quad=\quad 2 e^{\mathcal i 5\pi/3}\]

zepdrix (zepdrix):

z^7 shouldn't be too bad from there. Unless we're supposed to include all of the angles.. grr I always forget with these :p

OpenStudy (emilyjones284):

Wait im sorry im still confused because my teacher doesn't do it that way... Where would I go from there?

zepdrix (zepdrix):

From here,\[\Large\bf\sf z\quad=\quad 2\left(\frac{1}{2}-\frac{\sqrt3}{2}\mathcal i\right)\]Recognize that,\[\Large\bf\sf \cos\frac{5\pi}{3}\quad=\quad \frac{1}{2}\]\[\Large\bf\sf \sin\frac{5\pi}{3}\quad=\quad \frac{\sqrt3}{2}\]So we can write our expression as,\[\Large\bf\sf z\quad=\quad 2\left(\cos\frac{5\pi}{3}+\mathcal i \sin\frac{5\pi}{3}\right)\]That look familiar so far?

zepdrix (zepdrix):

Have you learned about `De Moivre's Theorem`?

zepdrix (zepdrix):

Woops sin(5pi/3) = -sqrt3/2. Typo there. Shouldn't change anything though.

OpenStudy (emilyjones284):

Oh okay yeah that's starting to look familiar

OpenStudy (emilyjones284):

So then where would you go from there to get z^7?

zepdrix (zepdrix):

So we have z. We'll raise both sides to the 7th power.

zepdrix (zepdrix):

The reason that's important is because you mustn't forget to put the power on the 2 as well.

zepdrix (zepdrix):

\[\Large\bf\sf z^7\quad=\quad 2^7\left(\cos\frac{5\pi}{3}+\mathcal i \sin\frac{5\pi}{3}\right)^7\]Have you learned about `De Moivre's Theorem`?

OpenStudy (emilyjones284):

Yeah we went over De Moivre's Theorem a little

zepdrix (zepdrix):

De Moivre's will allow us to bring the 7 inside of the argument of our sine and cosine.

zepdrix (zepdrix):

\[\Large\bf\sf z^7\quad=\quad 2^7\left(\cos\left[7\cdot\frac{5\pi}{3}\right]+\mathcal i \sin\left[7\cdot\frac{5\pi}{3}\right]\right)\]

OpenStudy (emilyjones284):

So 128=(cos 35pi/3 + isin 35pi/3)?

zepdrix (zepdrix):

Oh oh oh I screwed up on the angle I think. Pretty sure we want to use the `principle angle` which would be from -pi to pi. So for our angle we probably want to use -pi/3 instead of 5pi/3

zepdrix (zepdrix):

But uhhh yahhhh I think that puts us on the right track :U

OpenStudy (emilyjones284):

oh okay so 128= (cos 7pi/3+ isin 7pi/3)?

zepdrix (zepdrix):

Why is there an equals sign in there? +_+

OpenStudy (emilyjones284):

because 2^7 = 128

OpenStudy (emilyjones284):

oh oops nevermind haha I get it

zepdrix (zepdrix):

Ya something like that I think,\[\Large\bf\sf z^7\quad=\quad 128\left(\cos\left[-\frac{7\pi}{3}\right]+\mathcal i \sin\left[-\frac{7\pi}{3}\right]\right)\]negative on the angles. But then uhhh we can relize that umm....

zepdrix (zepdrix):

\[\Large\bf\sf -\frac{7\pi}{3}\quad=\quad -\frac{\pi}{3}-2\pi\]Since sine and cosine are periodic in 2pi, we can rewrite our angle as,\[\Large\bf\sf -\frac{\pi}{3}\]

zepdrix (zepdrix):

\[\Large\bf\sf z^7\quad=\quad 128\left(\cos\left[-\frac{\pi}{3}\right]+\mathcal i \sin\left[-\frac{\pi}{3}\right]\right)\]

zepdrix (zepdrix):

And then we can write our cosine and sine back in terms of their reference values. The same ones from before.\[\Large\bf\sf z^7\quad=\quad 128\left(\frac{1}{2}- \frac{\sqrt3}{2}\mathcal i\right)\]

zepdrix (zepdrix):

Then put a 2 back into the brackets.

zepdrix (zepdrix):

We kinda went in circles but it really just depends on what form you need your answer in.

OpenStudy (emilyjones284):

Oh so the answer would be \[128(1/2\left[ 2 \right]- \sqrt{3}/2i \left[ 2 \right])\]